VB code I have a code that tests a single textbox input making sure that the...

VB code

I have a code that tests a single textbox input making sure that the input is a number that is three long with no decimals, letters, or negative signs.

I'm trying to make a function that tests the three numbers to make sure that they are in decreasing order (543 or 951...) and a function that then reverses the three numbers (543 becomes 345)
Answer #1

VB function : use this function into your code:

Function fun(ByVal num As Integer) As Integer
Dim d(3) As Integer ' declaration of array for each digit*/
Dim i, j,temp As Integer
' seprating each digit*/
d(2)=num Mod 10
num=num / 10
d(1)= num Mod 10
d(0)=num / 10
' sorting digit */
For i = 0 To 3
For j = 2 To i + 1 Step -1
If (d(j) < d(j - 1)) Then
temp = d(j)
d(j) = d(j - 1)
d(j - 1) = temp
End If
Next
Next
' element of array to single number*/
num=(d(0)*100)+(d(1)*10)+(d(2))
return num
  

End Function
-----------------------------------------------------------------------------

IF YOU HAVE ANY QUERY PLEASE COMMENT DOWN BELOW

PLEASE GIVE A THUMBS UP